OpenStudy (anonymous):

has any one done Completing the Square for algrbr 2

OpenStudy (anonymous):

i have not sure if i remember it but can try and help

OpenStudy (anonymous):

did u do it

OpenStudy (anonymous):

do u go to online school

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

yes i have gone through alg 2

OpenStudy (anonymous):

Rewrite the equation in vertex form. Name the vertex and y-intercept. y = x2 – 10x + 15 y = (x – 10)2 –10 vertex: (–10, –10) y-intercept: (0, –10) y = (x – 5)2 – 10 vertex: (5, –10) y-intercept: (0, 15) y = (x – 5)2 + 40 vertex: (5, –10) y-intercept: (0, 15) y = (x – 10)2 + 20 vertex: (–10, –10) y-intercept: (0, –10)

OpenStudy (anonymous):

o Why not

OpenStudy (anonymous):

@Algebraic! could you give us a hand?

OpenStudy (anonymous):

dont remember how to do it

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

y = x2 – 10x + 15 y = (x – 10)2 –10

OpenStudy (anonymous):

so y = x^2-20x +90?

OpenStudy (anonymous):

y = x^2 – 10x + 15 y =x^2 -10x +25 -10 y =(x-5)(x-5) -10

OpenStudy (anonymous):

no y = x^2 – 10x + 15

OpenStudy (anonymous):

o never mind

OpenStudy (anonymous):

why did he leve

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

that's it...

OpenStudy (anonymous):

y =(x-5)^2 -10